OpenStudy (anonymous):

Speed of river current is 6mph.Ferry operator wants to buy new boat. Every day his route takes him 22.5 miles against the current and back to his dock and he needs to make this trip in a total of 9 hours. He has a boat in mind, but he can only test it on a lake where there is no current. How fast must the boat go on the lake in order for it to serve the ferry operator's needs?

OpenStudy (tkhunny):

Distance = Rate * Time Distance Up = Distance Down Rate Up = Boat Speed less River Current Rate Down = Boat Speed plus River Current Time Up + Time Down = 9 hours Go! Let's see what you get.

OpenStudy (anonymous):

is distance up half of 22.5 or is that the distance up and then the distance down

OpenStudy (tkhunny):

Whoops! Forgot to write that. Distance Up = Distance Down = 22.5 Miles

OpenStudy (anonymous):

so dist. up is 11.25 and down is 11.25?

OpenStudy (tkhunny):

Well, that's how I read it. "22.5 miles against the current ". If we don't get things as expected, we can decide that;s the WHOLE trip, rather than half the trip.

OpenStudy (tkhunny):

Distance Up = Distance Down = 22.5 Miles This is a 45 mile trip.

OpenStudy (anonymous):

that's how I am reading it that 22.5 is round trip

OpenStudy (anonymous):

but now u are saying round trip is 45 miles

OpenStudy (tkhunny):

Let's do it both ways and clear up the confusion. Are you solving with 22.5 for the whole trip and getting what an answer key says is "wrong".

OpenStudy (anonymous):

I don't have an answer key. I will work with round trip of 45 miles. do I know the boat speeed?

OpenStudy (tkhunny):

That's what we seek, isn't it?

OpenStudy (anonymous):

yes rate up 22.5=9x: x=2.5rate: 2.5mph-river current of 6 gives me a neative number. What am I doing wrong?

OpenStudy (anonymous):

rate down=2.5 +6=8mph

OpenStudy (tkhunny):

Don't be afraid to write things down. Trip Up 22.5 mi = (BoatSpeed - 6 mph)*(T Hours) Trip Down 22.5 mi = (BoatSpeed + 6 mph)*((9-T) Hours) How do you suggest we proceed given those clean and complete equations?

OpenStudy (tkhunny):

These equations should lead to a Quadratic Equation that can be factored. Boat Speed = 9 mph or Boat Speed = -4 mph. Get to that point. BTW: If we use 22.5 as the total round trip, we get Boat Speed = 7.3788 mph. I'm liking the other solution lots better.

OpenStudy (anonymous):

sorry, but I am not following. How can a boat speed be negative 4?

OpenStudy (tkhunny):

Excellent. You ARE following. That is EXACTLY the question I was hoping you would ask. Did you get the quadratic equation? Show me, please.